# Security Degradation in Iterative AI Code Generation: A Systematic Analysis of the Paradox

### v1

@@ Abstract
Index Terms:
Large Language Models, Security Vulnerabilities, AI-Generated Code, Iterative Feedback, Software Security, Secure Coding Practices, Feedback Loops, LLM Prompting Strategies
@@ IV-B Iteration-Specific Security Patterns
This pattern suggests a potential relationship between code modification cycles and security vulnerabilities. Statistical testing (repeated measures ANOVA) showed significant differences between early and late iterations (F(9,90) = 14.32, p ¡ 0.001, $\eta^{2}$ = 0.42), indicating a medium-to-large effect size.
@@ IV-C Vulnerability Type Analysis
Chi-square tests revealed significant differences in vulnerability type distributions across prompting strategies ($\chi^{2}$(33) = 172.4, p ¡ 0.001, Cramer’s V = 0.38). Efficiency-focused prompts were associated with memory safety issues (42.7%), while feature-focused prompts were associated with concurrency problems (30.4%).
@@ IV-D Code Evolution and Complexity
We found a positive correlation (r = 0.64, p ¡ 0.001) between code complexity increases and security vulnerability counts. For every 10% increase in complexity, we observed an average 14.3% increase in vulnerability count (95% CI: 10.7% - 17.9%). Multiple regression analysis controlling for prompting strategy and baseline code characteristics showed that complexity remained a significant predictor of vulnerability count ($\beta$ = 0.64, p ¡ 0.001, as detailed in Appendix B).
@@ B-A Repeated Measures ANOVA for Vulnerability Counts Across Iterations
Post-hoc Tukey HSD tests showed significant differences between iterations 1-3 and iterations 8-10 (p ¡ 0.001), but not between adjacent iterations.
@@ B-B Multiple Regression Analysis - Predicting Vulnerability Count
Iteration number 0.28 0.08 3.50 ¡0.001 [0.12, 0.44]
Model: $R^{2}$ = 0.67, F(5, 394) = 160.2, p ¡ 0.001

### v2

# Peer-reviewed and accepted in IEEE-ISTAS 2025 Security Degradation in Iterative AI Code Generation: A Systematic Analysis of the Paradox

@@ Abstract
PEER-REVIEWED AND ACCEPTED IN IEEE- ISTAS 2025 /https://attend.ieee.org/istas-2025/program/
@@ IV-B Iteration-Specific Security Patterns
This pattern suggests a potential relationship between code modification cycles and security vulnerabilities. Statistical testing (repeated measures ANOVA) showed significant differences between early and late iterations (F(9,90) = 14.32, p << 0.001, $\eta^{2}$ = 0.42), indicating a medium-to-large effect size.
@@ IV-C Vulnerability Type Analysis
Chi-square tests revealed significant differences in vulnerability type distributions across prompting strategies ($\chi^{2}$(33) = 172.4, p << 0.001, Cramer’s V = 0.38). Efficiency-focused prompts were associated with memory safety issues (42.7%), while feature-focused prompts were associated with concurrency problems (30.4%).
@@ IV-D Code Evolution and Complexity
We found a positive correlation (r = 0.64, p << 0.001) between code complexity increases and security vulnerability counts. For every 10% increase in complexity, we observed an average 14.3% increase in vulnerability count (95% CI: 10.7% - 17.9%). Multiple regression analysis controlling for prompting strategy and baseline code characteristics showed that complexity remained a significant predictor of vulnerability count ($\beta$ = 0.64, p << 0.001, as detailed in Appendix B).
@@ B-A Repeated Measures ANOVA for Vulnerability Counts Across Iterations
Post-hoc Tukey HSD tests showed significant differences between iterations 1-3 and iterations 8-10 (p << 0.001), but not between adjacent iterations.
@@ B-B Multiple Regression Analysis - Predicting Vulnerability Count
Iteration number 0.28 0.08 3.50 <<0.001 [0.12, 0.44]
Model: $R^{2}$ = 0.67, F(5, 394) = 160.2, p << 0.001